# Joint Processing and Transmission Energy Optimization for ISAC in Cell-Free Massive MIMO with URLLC

### v2

@@ Abstract
In this paper, we explore the concept of integrated sensing and communication (ISAC) within a downlink cell-free massive MIMO (multiple-input multiple-output) system featuring multi-static sensing and users requiring ultra-reliable low-latency communications (URLLC). Our work focuses on devising two algorithms that concurrently optimize power and blocklength allocation. The objectives of these algorithms are to minimize the energy consumption solely for transmission and the overall end-to-end (E2E) energy consumption, which encompasses sensing/communication processing and transmission. This is achieved while meeting the necessary requirements for sensing and URLLC. To address the inherent non-convexity of these optimization problems, we apply feasible point pursuit - successive convex approximation (FPP-SCA), concave-convex programming (CCP), and fractional programming. Our numerical analysis compares the performance of these algorithms in ISAC scenarios and against a baseline URLLC-only scenario without sensing integration. Results show the E2E energy minimization algorithm’s effectiveness, particularly in scenarios without sensing. Additionally, our study underscores the increasing prominence of energy consumption associated with sensing processing tasks as the number of sensing receive access points rises, indicating potential for energy savings by selectively deactivating APs. Furthermore, the results emphasize that although increasing the number of transmit APs can improve URLLC satisfaction and sensing update rate by reducing blocklength, careful management is essential to avoid excessive energy consumption and reduced detection probability.
@@ I Introduction
6G mobile networks are expected to offer various sensing-based applications such as autonomous vehicles, smart homes/cities/factories, remote healthcare, and location/environment-aware scenarios [1, 2]. Many of these applications, such as autonomous vehicles and industrial Internet-of-things (IIoT), demand ultra-reliable delivering of the sensing information to the user equipments (UEs) with a minimum reliability of 99.999% and within a low end-to-end (E2E) delay of less than 10-150 ms [3, 4, 5]. This demand necessitates two key aspects defined in 6G networks: integrated sensing and communication (ISAC) together with ultra-reliable low-latency communication (URLLC). For most URLLC applications, short codewords are usually needed to satisfy latency constraints where codes with short blocklengths, e.g., 50-400 symbols are employed. Short blocklengths result in decoding errors; therefore, to ensure reliability, it is essential to consider the finite blocklength regime for modeling the decoding error probability (DEP) [6, 7, 8].
@@
In mission-critical applications such as traffic control and autonomous vehicles, real-time transmission of sensing information to URLLC UEs is essential, as the speed of information updates directly impacts the effectiveness of these applications. In this paper, we introduce the concept of the “sensing update rate (SUR),” which quantifies the number of sensing measurements and updates per time unit [9]. The SUR is influenced by the blocklength, with longer blocklengths resulting in lower update rates due to the extended duration of each sensing operation. This dependency creates an opportunity to optimize blocklength to balance reliable communication with high SUR. Although URLLC and ISAC have typically been studied as separate domains, a joint design and optimization approach is required to meet the distinct requirements of both and maximize integration gains.
@@
Cell-free massive multiple-input multiple-output (CF-mMIMO) has been shown to be a good candidate to facilitate stringent requirements of URLLC applications [8, 10, 11] as it provides high reliability by handling large path-loss variations and improves cell-edge user performance. CF-mMIMO can also facilitate ISAC implementations leveraging distributed access points (APs) to deploy bi-static/ multi-static sensing eliminating the need for full-duplex capability at the APs [12, 13, 14, 15]. The central processing of network functions within a centralized radio access network (C-RAN) architecture facilitates coordination among distributed APs, enabling joint transmission and processing in CF-mMIMO[16]. C-RAN has evolved into virtualized C-RAN and, with the open interfaces, further into Open RAN (O-RAN), where virtualization enables the use of general-purpose servers for jointly processing of baseband and sensing signals.
However, deploying these technologies in 6G is anticipated to result in substantial energy consumption. In CF-mMIMO networks, baseband processing combined with sensing tasks is expected to lead to higher energy usage in ISAC networks compared to conventional communication systems. The impact of sensing on energy consumption in such networks has not been well-studied. Existing studies on green CF-mMIMO primarily focus on optimizing transmit power consumption in radio access, with limited work addressing processing energy, aside from [17]. Virtualized C-RAN architecture facilitates centralized processing and enables the sharing of processing resources between APs and tasks, paving the way for end-to-end (E2E) optimization of both processing and radio resources in ISAC networks. This motivates a comprehensive study of the E2E energy consumption in ISAC networks, considering not only energy consumption for transmission but also for processing tasks. The primary research question addressed in this paper is how to incorporate sensing in CF-mMIMO systems with URLLC while minimizing E2E energy consumption.
@@ I-A Literature Review
The finite blocklength regime and URLLC have primarily been studied in the context of cellular networks (see [18, 19, 20] and references therein), with limited research addressing URLLC in CF-mMIMO networks. The potentials of massive MIMO in meeting URLLC requirements for industrial automation are investigated in [21, 22, 23]. Network availability is investigated in [21] and [22], and [23] proposes a power allocation algorithm to maximize the minimum signal-to-interference-plus-noise ratio (SINR). The authors in [24] propose two power optimization algorithms to provide URLLC for traditional ground UEs and unmanned aerial vehicles (UAVs). The objectives in [24] are maximizing either sum or minimum URLLC rate and they study neither energy minimization nor blocklength optimization. Max-min rate optimization is also studied in [11], where the authors adapt a special class of the conjugate beamforming for a CF-mMIMO with single-antenna APs. Moreover, the authors aim to address an energy efficiency optimization problem.
@@
Joint power and blocklength optimization is studied in [25] and [26] to optimize the reliability in a factory automation scenario and heterogeneous cellular networks, respectively. Cell-free networks as well as energy minimization are not studied in these works. Moreover, sensing requirements are not considered in [23, 21, 24, 11, 25, 26].
@@
There are few works that jointly consider URLLC and ISAC. In [5], a joint precoding scheme is proposed to minimize transmit power, satisfying sensing and delay requirements. Moreover, joint ISAC beamforming and scheduling design is addressed in [27] for periodic and aperiodic traffic.
@@
The consideration of E2E energy-awareness has been explored in various contexts, as reflected in prior works such as [28, 29, 17, 30, 31]. In particular, [30] studied fully virtualized E2E power minimization problem for CF-mMIMO on O-RAN architecture by taking the radio, fronthaul, and processing resources into account. Joint user scheduling and power allocation scheme for cell-free ISAC systems is studied in [32] that aims to maximize the sum-rate of the communication UEs and the sensing targets. However, the authors do not consider either URLLC requirements or energy minimization.
@@
To the best of our knowledge, for the first time, ISAC in CF-mMIMO with URLLC UEs and with the focus on energy efficiency was studied in our previous work[33]. In [33], we propose a successive convex approximation-based power allocation algorithm that maximizes energy efficiency while satisfying the sensing and URLLC requirements for a given blocklength. However, the impact of blocklength optimization is not considered in [33]. In this paper, we extend [33] from an E2E energy minimization perspective and aim to address a joint blocklength and power optimization problem. Here, E2E refers to considering energy consumption at both radio site and the central cloud.
@@ I-B Contributions
In this paper, we study E2E energy consumption in a CF-mMIMO system with URLLC UEs and multi-static sensing in a cluttered environment. We sense one candidate location of the target utilizing communication signals and an additional sensing signal. The sensing signals can cause interference for the UEs. Thus, we design the sensing precoding vectors to null the interference for the UEs. We consider a maximum DEP threshold, representing the reliability requirement, together with a maximum transmission delay threshold as the URLLC requirements and a minimum SINR as the sensing requirement. The latter is motivated by the fact that given a fixed false alarm probability, detection probability increases with a higher sensing SINR [34], [9, Chap. 3 and 15].
Different from [33], we conduct a comprehensive analysis of the processing requirements, specifically in terms of giga-operations per second (GOPS), and examine its impact on processing energy consumption. The main contributions of this paper are outlined as follows:
  • We propose two joint blocklength and power allocation algorithms aimed at minimizing transmission energy consumption and E2E energy consumption, respectively, while satisfying URLLC and sensing requirements. To address the non-convex optimization challenges, we employ feasible point pursuit - successive convex approximation (FPP-SCA), concave-convex programming (CCP), and fractional programming.
  • We analyze the impact of E2E energy minimization, covering both radio and processing energy consumption, within the proposed ISAC scenario and a comparative scenario that considers only URLLC requirements.
  • We examine two distinct target detection schemes, comparing their performance in terms of E2E energy consumption, detection probability, and sensing update rate.
  • We conduct sensitivity analyses to evaluate the effects of DEP threshold, sensing SINR threshold, and the number of ISAC transmit APs and sensing receive APs on E2E energy consumption, detection probability, and sensing update rate.
The rest of the paper is organized as follows: Section II introduces the system model. Section III provides URLLC analysis, considering DEP and delay in finite blocklength regime. Section IV covers the sensing analysis. Section V describes the E2E power model and derives GOPS analysis for both communication and sensing. Optimization problems are presented in Section VI, followed by numerical results and conclusions in Sections VII and VIII, respectively.
@@ II System Model
We study ISAC in a CF-mMIMO system in URLLC scenarios. The system adopts a virtualized C-RAN architecture [35] for uplink channel estimation and downlink communication, as well as multi-static sensing as shown in Fig. 1. All the APs are interconnected via fronthaul links to the edge cloud and operate in full synchronization. We consider the original form of CF-mMIMO [8], wherein all the $N_{\rm tx}$ ISAC APs jointly serve the $N_{\rm ue}$ URLLC UEs by transmitting precoded signals containing both communication and sensing symbols.
Figure 1: Illustration of an ISAC system in CF-mMIMO with URLLC UEs.
@@
II-A Downlink Communication
Let $s_{i}[m]$ and $s_{0}[m]$ represent the downlink communication symbol for UE $i$ and sensing symbol, respectively at time instance $m$. The symbols are independent and have zero mean and unit power. Moreover, let $\rho_{i}\geq 0$ and $\rho_{0}\geq 0$ be, respectively, the power control coefficients for UE $i$ and the target, which are fixed throughout the transmission. Then, the transmitted signal from transmit AP $k\in\{1,\cdots,N_{\rm tx}\}$ at time instance $m$ is
$\vspace{-1mm}\textbf{x}_{k}[m]=\sum_{i=0}^{N_{\rm ue}}\sqrt{\rho_{i}}\textbf{w}_{i,k}s_{i}[m]=\textbf{W}_{k}\textbf{D}_{\rm s}[m]\bm{\rho},\vspace{-1mm}$ (1)
@@ II-A Downlink Communication
The communication channels are modeled as spatially correlated Rician fading which are assumed to remain constant during each coherence block, and the channel realizations are independent of each other. Let $\textbf{h}_{i,k}\in\mathbb{C}^{M}$ denote the channel between ISAC AP $k$ and UE $i$, modeled as
@@
which consists of a semi-deterministic line-of-sight (LOS) path, represented by $e^{j\varphi_{i,k}}\bar{\textbf{h}}_{i,k}$ with unknown phase-shift $\varphi_{i,k}\sim\mathcal{U}[0,2\pi)$, i.e., uniformly distributed on $[0,2\pi)$, and a stochastic non-LOS (NLOS) component $\tilde{\textbf{h}}_{i,k}\sim\mathcal{CN}(\textbf{0},\textbf{R}_{i,k})$ with the spatial correlation matrix $\textbf{R}_{i,k}\in\mathbb{C}^{M\times M}$. Both $\bar{\textbf{h}}_{i,k}$ and $\textbf{R}_{i,k}$ include the combined effect of geometric path loss and shadowing. We concatenate the channel vector $\textbf{h}_{i,k}$ in the collective channel vector
@@
where $n_{i}[m]\!\sim\!\mathcal{CN}(0,\sigma_{n}^{2})$ is the independent receiver noise at UE $i$ at time instance $m$ and the collective precoding vectors
${\textbf{w}_{i}=\left[\textbf{w}_{i,1}^{T}\ \textbf{w}_{i,2}^{T}\ \ldots\ \textbf{w}_{i,N_{\rm tx}}^{T}\right]^{T}\in\mathbb{C}^{N_{\rm tx}M}},$ (5)
for $i=1,\ldots,N_{\rm ue}$, and
$\textbf{w}_{0}=\left[\textbf{w}_{0,1}^{T}\ \textbf{w}_{0,2}^{T}\ \ldots\ \textbf{w}_{0,N_{\rm tx}}^{T}\right]^{T}\in\mathbb{C}^{N_{\rm tx}M},$ (6)
are the centralized precoding vectors.
Using the independence of the data and sensing signals, the average transmit power for transmit AP $k\in\{1,\ldots,N_{\rm tx}\}$ is
$P_{k}=\sum_{i=0}^{N_{\rm ue}}\rho_{i}\mathbb{E}\left\{\|\textbf{w}_{i,k}\|^{2}\right\},\quad k=1,\ldots,N_{\rm tx}$ (7)
which should not exceed the maximum power limit $P_{\rm tx}$.
In the subsequent section, we will present the specifics of the linear minimum mean-squared error (LMMSE)-based channel estimation.
II-B LMMSE-Based Channel Estimation
Within the URLLC transmission time comprising $L$ symbols, $L_{p}<L$ symbols are allocated for pilot transmission. A collection of $L_{p}$ mutually orthogonal pilot sequences is represented by $\bm{\phi}_{1},\ldots,\bm{\phi}_{L_{p}}\in\mathbb{C}^{L_{p}}$, where $\|\bm{\phi}_{t}\|^{2}=L_{p}$ for $t=1,\ldots,L_{p}$. These pilot sequences are assigned to the $N_{\rm ue}$ UEs. If the number of pilot sequences is less than the number of UEs (i.e., $L_{p}<N_{\rm ue}$), then each pilot sequence may be assigned to multiple UEs. Let $t_{i}\in\{1,\ldots,L_{p}\}$ denote the index of the pilot sequence that is assigned to UE $i$ and we define $\mathcal{P}_{i}=\{j\in\{1,\ldots,N_{\rm ue}\}|t_{j}=t_{i}\}$ as the set of UEs that share the same pilot with UE $i$.
The received signal at the ISAC AP $k$ during the entire pilot transmission is denoted by $\textbf{Y}_{k}^{p}\in\mathbb{C}^{M\times L_{p}}$ where
$\textbf{Y}_{k}^{p}=\sum_{i=1}^{N_{\rm ue}}\sqrt{\eta_{i}}\textbf{h}_{i,k}\bm{\phi}_{t_{i}}^{T}+\textbf{N}_{k},$ (8)
where $\eta_{i}>0$ is the transmit power by UE $i$ and $\textbf{N}_{k}\in\mathbb{C}^{M\times L_{p}}$ is the received noise at AP $k$ with independent and identically distributed (i.i.d.) $\mathcal{CN}(0,\sigma_{n}^{2})$ entries. We assume that $\bar{\textbf{h}}_{i,k}$ and $\textbf{R}_{i,k}$, representing the long-term statistics for the channel in (2), are known while the random phase-shift $\varphi_{i,k}$ is unknown. With respect to estimating $\textbf{h}_{i,k}$, let us form $\textbf{y}_{t_{i},k}^{p}=\textbf{Y}_{k}^{p}\bm{\phi}_{t_{i}}^{*}\in\mathbb{C}^{M}$, where
$\textbf{y}_{t_{i},k}^{p}=\underbrace{\sqrt{\eta_{i}}\,L_{p}\textbf{h}_{i,k}}_{\textrm{Desired part}}+\underbrace{\sum_{j\in\mathcal{P}_{i}\setminus\{i\}}\sqrt{\eta_{j}}\,L_{p}\textbf{h}_{j,k}}_{\textrm{Interference part}}+\underbrace{\textbf{n}_{t_{i},k}}_{\textrm{Noise}},$ (9)
@@ II-B LMMSE-Based Channel Estimation
where $\textbf{n}_{t_{i},k}=\textbf{N}_{k}\bm{\phi}_{t_{i}}^{*}\sim\mathcal{CN}(\textbf{0},\sigma_{n}^{2}L_{p}\textbf{I}_{M})$. If the number of UEs is greater than the number of mutually orthogonal pilot sequences, then each pilot sequence may be assigned to multiple UEs using the pilot assignment algorithm in [35, Alg. 4.1]. The LMMSE estimate of $\textbf{h}_{i,k}$ is given by [37]
$\hat{\textbf{h}}_{i,k}=\sqrt{\eta_{i}}\textbf{R}_{i,k}^{\prime}\left(\bm{\Psi}^{\prime}_{t_{i},k}\right)^{-1}\textbf{y}_{t_{i},k}^{p},$ (10)
where $\textbf{R}_{i,k}^{\prime}$ and $\bm{\Psi}_{t_{i},k}^{\prime}$ are the spatial correlation matrix corresponding to channel $\textbf{h}_{i,k}$ and the correlation matrix of the received signal, respectively, given by
$\textbf{R}_{i,k}^{\prime}\triangleq\textbf{R}_{i,k}+\bar{\textbf{h}}_{i,k}\bar{\textbf{h}}_{i,k}^{H}$ (11)
$\bm{\Psi}^{\prime}_{t_{i},k}\triangleq\sum_{j\in\mathcal{P}_{i}}\eta_{j}L_{p}\textbf{R}_{j,k}^{\prime}+\sigma_{n}^{2}\textbf{I}_{M}.$ (12)
It is worth noting that deriving MMSE channel estimation proves challenging due to the non-Gaussian distribution of channels resulting from random phase-shifts in the LOS paths. For the sake of tractability, we adhere to the use of LMMSE-based channel estimation in this paper.
II-C Multi-Static Sensing
We investigate multi-static sensing, involving multiple transmit and receive APs in the network. The network engages in target location sensing during the downlink phase. We assume the existence of a LOS connection between the target location and each transmit/receive AP.

### v3

# End-to-End Energy Saving in Cell-Free Massive MIMO ISAC for Ultra-Reliable Target-Aware Actuation

@@ Abstract
[…]
@@ I Introduction
6G mobile networks are expected to offer various sensing-based applications such as autonomous vehicles, smart homes/cities/factories, remote healthcare, industrial Internet-of-things (IIoT), and robot control in target/environment-aware scenarios [1, 2]. These applications introduce ultra-reliable target-aware actuation use cases, which involve scenarios where a communication and sensing system must detect the presence, position, and motion of a target (e.g., a human or object) with high accuracy and minimal delay, and then trigger an actuation response (e.g., controlling a robot, vehicle, or machine) with ultra-reliability and low latency.
To support such use cases, the system must deliver sensing information to user equipment (UE) with a minimum reliability of 99.999% and an end-to-end (E2E) latency of less than 10–150 ms [3, 4, 5]. This demand necessitates two key aspects defined in 6G networks: integrated sensing and communication (ISAC) together with ultra-reliable low-latency communication (URLLC).
@@
In ultra-reliable target-aware actuation use cases such as traffic control and autonomous vehicles, real-time transmission of sensing information to URLLC UEs is essential, as the timeliness of updates directly affects the performance and safety. To capture the frequency at which sensing information must be updated, this paper introduces the concept of the “refreshing rate”, defined by 3GPP [6] as the number of sensing measurements and updates delivered per unit time [7]. The refreshing rate is closely tied to the communication blocklength: longer blocklengths lead to lower refreshing rates due to the increased time required for each sensing operation. This trade-off highlights the potential for optimizing blocklength to achieve a balance between reliable communication and a high refreshing rate. Although URLLC and ISAC have typically been studied as separate domains, a joint design and optimization approach is required to meet the distinct requirements of both and maximize integration gains.
@@
Cell-free massive multiple-input multiple-output (CF-mMIMO) has emerged as a strong candidate for meeting the stringent requirements of URLLC applications [8, 9, 10]. It offers high reliability by compensating for large path-loss variations and significantly improves the performance of cell-edge UEs. Moreover, CF-mMIMO is well-suited for implementing ISAC, as its distributed access points (APs) can support bi-static and multi-static sensing configurations [11, 12, 13, 14]. These configurations eliminate the need for full-duplex capabilities at individual APs but require effective coordination among them. Centralized radio access network (C-RAN) architectures, which support centralized processing of network functions, enables joint transmission/reception and processing [15]. This centralized approach is also essential for facilitating bi-static and multi-static sensing.
While CF-mMIMO networks offer significant performance advantages, the integration of sensing functionalities is anticipated to substantially increase transmission energy consumption, resulting in up to a tenfold reduction in energy efficiency [16]. This challenge is further compounded by the fact that the combined demands of baseband processing and sensing operations may result in higher processing energy consumption compared to conventional communication-only systems. Most existing research on green CF-mMIMO without sensing integration has primarily focused on optimizing transmit power, with relatively limited attention given to processing energy consumption, aside from a few notable exceptions [17, 18]. Moreover, the impact of sensing functionalities on processing energy consumption in such networks remains insufficiently explored.
It is important to mention that communication system efficiency is typically measured in terms of data rate while sensing tasks are evaluated by task-specific metrics such as target detection success rates. The execution of a sensing task involves not only signal transmission and reception of reflections from the target but also extensive processing to extract the desired information. This post-reception processing is a fundamental component of completing a sensing task, and thus must be explicitly accounted for in system design and analysis. This distinction highlights the need for a comprehensive E2E energy consumption analysis that accounts for both transmission and computation costs.
The virtualization capabilities offered by C-RAN architectures facilitate centralized resource management and flexible allocation of processing resources across APs and tasks. This architectural flexibility presents a promising opportunity for E2E optimization of both processing and radio resources in ISAC networks.
Motivated by these considerations, this paper addresses the following key research question: How can sensing be effectively integrated into CF-mMIMO systems with CRAN architecture and supporting URLLC, while minimizing E2E energy consumption?
@@ I-A Literature Review
The finite blocklength regime and URLLC have primarily been studied in the context of cellular networks (see [19, 20, 21] and references therein), with limited research addressing URLLC in CF-mMIMO networks. The potentials of massive MIMO in meeting URLLC requirements for industrial automation are investigated in [22, 23, 24]. Network availability is investigated in [22] and [23], and [24] proposes a power allocation algorithm to maximize the minimum signal-to-interference-plus-noise ratio (SINR). The authors in [25] propose two power optimization algorithms to provide URLLC for traditional ground UEs and unmanned aerial vehicles (UAVs) in a CF-mMIMO system. The objectives in [25] are maximizing either the sum or the minimum URLLC rate, and they study neither energy minimization nor blocklength optimization. Max-min rate optimization is also studied in [10], where the authors adapt a special class of conjugate beamforming for a CF-mMIMO with single-antenna APs. Moreover, the authors aim to address an energy efficiency optimization problem.
@@
Joint power and blocklength optimization is studied in [26] and [27] to optimize the reliability in a factory automation scenario and heterogeneous cellular networks, respectively. Cell-free networks as well as energy minimization are not studied in these works. Moreover, sensing requirements are not considered in [24, 22, 25, 10, 26, 27, 23].
@@
There are few works that jointly consider URLLC and ISAC. In [5], a joint precoding scheme is proposed to minimize transmit power, satisfying sensing and delay requirements. Moreover, joint ISAC beamforming and scheduling design is addressed in [28] and [29] with a focus on the coexistence of periodic and aperiodic traffic to balance the tradeoff between their corresponding performance. The aperiodic traffic is triggered by sensing information about the stochastic environment.
@@
The consideration of E2E energy-awareness has been explored in various contexts, as reflected in prior works such as [30, 31, 17, 18, 32]. In particular, [18] studied fully virtualized E2E power minimization problem for CF-mMIMO on O-RAN architecture by taking the radio, fronthaul, and processing resources into account. Joint UE scheduling and power allocation scheme for cell-free ISAC systems is studied in [33], which aims to maximize the sum rate of the communication UEs and the sensing targets. However, the authors do not consider either URLLC requirements or energy minimization.
TABLE I: NOVELTY COMPARISON OF OUR PAPER TO THE EXISTING WORKS IN LITERATURE
[19, 20, 21] [22, 23, 24] [25] [10] [26, 27] [5] [28, 29] [30, 31, 17, 18, 32] [16] This paper
ISAC
URLLC
MIMO
Cell-free networks
Multi-static sensing
Target detection
E2E energy consumption
Blocklength optimization
@@
Table I presents a comparison highlighting the novelty of this work relative to existing studies. To the best of our knowledge, our previous work [16] was the first to investigate ISAC in CF-mMIMO systems with URLLC UEs, with a specific focus on energy efficiency. However, the impact of blocklength optimization was not addressed in that study. In this paper, we formulate and solve a joint blocklength and power optimization problem from an E2E energy minimization perspective. E2E energy consumption encompasses energy used both at the distributed radio APs and at the centralized cloud infrastructure.
Figure 1: ISAC network architecture in CF-mMIMO with URLLC UEs and a joint resource allocation mechanism.
@@ I-B Contributions
Integration of communication and sensing functionalities in a CF-mMIMO network is depicted in Fig. 1. The objective is to jointly optimize blocklength and power allocation to minimize E2E energy consumption across both the radio access and edge cloud domains. URLLC requirements—such as reliability and latency—and sensing requirements—such as refreshing rate and accuracy—are jointly addressed through coordinated resource allocation. The edge cloud manages centralized processing for both sensing and communication, while distributed APs enable multi-static sensing alongside communication services. This architecture leverages shared radio and computation resources to facilitate joint performance optimization across sensing and communication domains.
In this paper, we study E2E energy consumption in a CF-mMIMO system with URLLC UEs and multi-static sensing in a cluttered environment. We sense a candidate location of the potential target utilizing communication signals and an additional sensing signal. Since the sensing signals can cause interference for the UEs, we employ null-space precoding vector for the sensing signal to null the interference to the UEs. The main contributions of this paper are outlined as follows:
  • We conduct a comprehensive analysis of the processing requirements, specifically in terms of giga-operations per second (GOPS) for communication and sensing tasks at the cloud, and provide analytical expression of the cloud energy consumption.
  • We present an E2E energy minimization algorithm by jointly optimizing the blocklength and power and compare it with two baseline algorithms. The first baseline algorithm aims to minimize the E2E energy consumption for only CF-mMIMO systems with URLLC without sensing capability. The second baseline considers sensing in the system while it aims to minimize only transmission energy consumption at the radio sites. To address the non-convex optimization challenges, we employ feasible point pursuit - successive convex approximation (FPP-SCA), concave-convex programming (CCP), and fractional programming.
  • We evaluate two distinct target detection schemes based on the maximum a posteriori ratio test (MAPRT), each with different processing requirements—namely, the clutter-unaware detector and the clutter-aware detector. Their performance is compared in terms of E2E energy consumption and detection probability.
  • We conduct sensitivity analyses to evaluate the effects of communication and sensing requirements, number of antennas per AP, and sensing receive APs on E2E energy consumption and sensing performance.
The rest of the paper is organized as follows: Section II introduces the system model. Section III provides URLLC analysis, considering decoding error probability (DEP) and delay in the finite blocklength regime. Section IV covers the sensing analysis. Section V describes the E2E power model and derives GOPS analysis for both communication and sensing. Optimization problems are presented in Section VI, followed by numerical results and conclusions in Sections VII and VIII, respectively.
[…]
@@ II System Model
We study ISAC in a CF-mMIMO system in URLLC scenarios. The system adopts a virtualized C-RAN architecture [34] for uplink channel estimation and downlink communication, as well as multi-static sensing as shown in Fig. 2. All the APs are interconnected via fronthaul links to the edge cloud and operate in full synchronization. Notice that synchronization presents a major challenge in multi-static sensing, as asynchronous transmitters and receivers can lead to ambiguity and performance degradation. While this issue has been studied in the context of communication in cell-free networks [35], the synchronization mechanisms already developed for CF-mMIMO communication make it a promising infrastructure for implementing advanced sensing techniques. This motivates our focus on exploring other critical aspects of CF-mMIMO-based ISAC systems. Another important consideration is the significant fronthaul capacity requirement imposed by centralized signal processing for both communication and sensing functions [36]. While this paper adopts a centralized processing framework to evaluate the full potential of CF-mMIMO ISAC systems, addressing fronthaul limitations remains an open research direction.
We consider the original form of CF-mMIMO [8], wherein all the $N_{\rm tx}$ ISAC APs jointly serve the $N_{\rm ue}$ URLLC UEs by transmitting precoded signals containing both communication and sensing symbols.
Figure 2: ISAC system model in CF-mMIMO with URLLC UEs.
@@
II-A Downlink ISAC Transmission
Let $s_{i}[m]$ and $s_{0}[m]$ represent the downlink communication symbol for UE $i$ and sensing symbol, respectively at time instance $m$. The symbols are independent and have zero mean and unit power. Moreover, let $\rho_{i}\geq 0$ and $\rho_{0}\geq 0$ be, respectively, the power control coefficients for UE $i$ and the target, which are fixed throughout the transmission. Then, the transmitted signal from transmit AP $k\in\{1,\ldots,N_{\rm tx}\}$ at time instance $m$ is
$\vskip-2.84526pt\textbf{x}_{k}[m]=\sum_{i=0}^{N_{\rm ue}}\sqrt{\rho_{i}}\textbf{w}_{i,k}s_{i}[m]=\textbf{W}_{k}\textbf{D}_{\rm s}[m]\bm{\rho},\vskip-2.84526pt$ (1)
@@ II-A Downlink ISAC Transmission
The communication channels are modeled as spatially correlated Rician fading, which are assumed to remain constant during each coherence block, and the channel realizations are independent of each other. Let $\textbf{h}_{i,k}\in\mathbb{C}^{M}$ denote the channel between ISAC AP $k$ and UE $i$, modeled as
@@
which consists of a semi-deterministic line-of-sight (LOS) path, represented by $e^{j\varphi_{i,k}}\bar{\textbf{h}}_{i,k}$ with unknown phase-shift $\varphi_{i,k}\sim\mathcal{U}[0,2\pi)$, i.e., uniformly distributed on $[0,2\pi)$, and a stochastic non-LOS (NLOS) component $\tilde{\textbf{h}}_{i,k}\sim\mathcal{CN}(\textbf{0},\textbf{R}_{i,k})$ with the spatial correlation matrix $\textbf{R}_{i,k}\in\mathbb{C}^{M\times M}$. Both $\bar{\textbf{h}}_{i,k}$ and $\textbf{R}_{i,k}$ include the combined effect of geometric path loss and shadowing. We concatenate the channel vectors $\textbf{h}_{i,k}$ in the collective channel vector
@@
where $n_{i}[m]\!\sim\!\mathcal{CN}(0,\sigma_{n}^{2})$ is the independent receiver noise at UE $i$ at time instance $m$ and the collective precoding vectors $\textbf{w}_{i}=\left[\textbf{w}_{i,1}^{T}\ \textbf{w}_{i,2}^{T}\ \ldots\ \textbf{w}_{i,N_{\rm tx}}^{T}\right]^{T}\in\mathbb{C}^{N_{\rm tx}M},$ for $i=1,\ldots,N_{\rm ue}$, and $\textbf{w}_{0}=\left[\textbf{w}_{0,1}^{T}\ \textbf{w}_{0,2}^{T}\ \ldots\ \textbf{w}_{0,N_{\rm tx}}^{T}\right]^{T}\in\mathbb{C}^{N_{\rm tx}M}$ are the centralized precoding vectors.
@@
where $\delta$ is the regularization parameter, and $\hat{\textbf{h}}_{j}=\begin{bmatrix}\hat{\textbf{h}}_{j,1}^{T}&\ldots&\hat{\textbf{h}}_{j,N_{\rm tx}}^{T}\end{bmatrix}^{T}\in\mathbb{C}^{N_{\rm tx}M}$ is the linear minimum mean-squared error (LMMSE) channel estimate of the communication channel $\textbf{h}_{j}$, obtained as in [38]111 We omit the explanations from this paper due to the space limitation.. If the number of UEs is larger than the number of mutually orthogonal pilot sequences, then each pilot sequence may be assigned to multiple UEs using the pilot assignment algorithm in [34, Algorithm 4.1].
and U is the unitary matrix with the orthogonal columns that span the column space of the matrix $\begin{bmatrix}\hat{\textbf{h}}_{1}&\ldots&\hat{\textbf{h}}_{N_{\rm ue}}\end{bmatrix}$. $\textbf{h}_{0}=\begin{bmatrix}\sqrt{\beta_{1}}\textbf{a}^{T}(\varphi_{1},\vartheta_{1})&\ldots&\sqrt{\beta_{N_{\rm tx}}}\textbf{a}^{T}(\varphi_{N_{\rm tx}},\vartheta_{N_{\rm tx}})\end{bmatrix}^{T}\in\mathbb{C}^{N_{\rm tx}M}$ is the concatenated sensing channel between all the ISAC APs and the target, including the corresponding channel gains $\beta_{k}$ and the array response vectors $\textbf{a}(\varphi_{k},\vartheta_{k})$ for $k=1,\ldots,N_{\rm tx}$.
II-B Multi-Static Sensing